# noitemsep,topsep=0pt,parsep=0pt,partopsep=0pt

### v2

@@ Abstract
DistServe assigns prefill and decoding computation to different GPUs, hence eliminating prefill-decoding interferences. Given the application’s TTFT and TPOT requirements, DistServe co-optimizes the resource allocation and parallelism strategy tailored for each phase. DistServe also places the two phases according to the serving cluster’s bandwidth to minimize the communication caused by disaggregation. As a result, DistServe significantly improves LLM serving performance in terms of the maximum rate that can be served within both TTFT and TPOT constraints on each GPU. Our evaluations show that on various popular LLMs, applications, and latency requirements, DistServe can serve 4.48$\times$ more requests or 10.2$\times$ tighter SLO, compared to state-of-the-art systems, while staying within latency constraints for $>90\%$ of requests.
@@ 1 Introduction
Based on the above insights, in this work, we build DistServe, a goodput-optimized LLM serving system by disaggregating the prefill and decoding phases. Given TTFT and TPOT requirements, DistServe first scales each phase independently by co-optimizing the GPU allocation and parallelism strategies of the prefill and decoding phase assuming serving a single model replica. The optimization ensures maximizing the per-GPU goodput and may assign different numbers of GPUs and parallelism strategies to each phase depending on their respective latency requirements. DistServe then scales this allocation to multiple instances via replication until meeting the user-required traffic rate (§4). DistServe also features an algorithm to place the prefill and decoding computation according to their allocation schemes and the cluster’s bandwidth to minimize the overhead of communicating intermediate states between phases.
We implement DistServe as an orchestration layer on top of the LLM inference engine. We evaluate DistServe on various LLMs, varying the workloads based on three important real-world LLM applications: chatbots, programming assistant, and document summary. Compared to state-of-the-art solutions, DistServe can serve up to 4.48×4.48$\times$4.48 × more requests under latency constraints. Our contributions are:
  • Identify the problems of prefill-decoding interference and resource coupling in existing LLM serving systems and propose to disaggregate the prefill and decoding phases.
  • Design a novel placement algorithm to automatically choose the goodput-optimal schema for prefill and decoding instances.
@@ 2.1 LLM Inference
The prefill step deals with a new sequence, often comprising many tokens, and processes these tokens concurrently. Unlike prefill, each decoding step only processes one new token generated by the previous step. This leads to significant computational differences between the two phases. When dealing with user prompts that are not brief, the prefill step tends to be computation-bound. For instance, for a 13B LLM, computing the prefill of a 512-token sequence makes an A100 compute-bound. The larger the model, the shorter sequence is needed to turn the prefill step compute-bound (see §3.1). In contrast, the decoding phase, despite processing only one new token per step, incurs a similar level of I/O to the prefill phase, making it constrained by the GPU’s memory bandwidth.
(a) Input length = 128            (b) Input length = 1024
Figure 2: Execution time for one batch when serving an LLM with 13B parameters as batch size increases. Compared between a decoding-only batch and a batch adding just one prefill request.
@@ 2.2 LLM Serving Optimization
Batching. Current serving systems [45, 28, 8] utilize a batching technique known as continuous batching. This method batches the prefill of new requests with the decoding of ongoing ones. It boosts the GPU utilization and maximizes the overall system throughput – tokens generated per second across all users and requests. However, as mentioned in §1 and elaborated later in §2.3, this approach leads to trade-offs between TTFT and TPOT. An advanced variant of continuous batching [8] attempts to balance TTFT and TPOT by segmenting prefill and attaching decoding jobs in a manner that avoids exceeding GPU performance limits – but essentially, it trades TTFT for TPOT. In summary, batching prefill and decoding invariably leads to compromises in either TTFT or TPOT.
Model parallelism. In LLM serving, model parallelism is generally divided as intra- and inter-operator parallelisms [29, 50, 39]. Both can be used to support larger models but may impact serving performance differently. Intra-operator parallelism partitions computationally intensive operators, such as matrix multiplications, across multiple GPUs, accelerating computation but causing substantial communication. It reduces the execution time222we emphasize “execution time” instead of latency here because latency comprises both execution time and queuing delay., hence latency, particularly for TTFT of the prefill phase, but requires high bandwidth connectivity between GPUs (e.g., NVLink). Inter-operator parallelism organizes LLM layers into stages, each running on a GPU to form pipelines. It moderately increases execution time due to inter-stage communication, but linearly scales the system’s rate capacity with each added GPU. In this paper, we reveal an additional benefit of model parallelism: reduced queuing delay of both prefill and decoding phases, steaming from shorter execution time. We delve into this further in §3. Besides model parallelism, replicating a model instance, irrespective of its model parallelism configurations, linearly scales the system’s rate capacity.
@@ 2.3 Problems and Opportunities
Resource and parallelism coupling. Colocating prefill and decoding phases on the same GPUs unavoidably share their resource and parallelism settings. However, each phase has its unique computational characteristic and latency requirement that calls for more heterogeneous resource allocation. For example, the prefill phase benefits from more GPUs and intra-op parallelism to reduce execution time to meet the tight SLO on TTFT. The decoding phase can handle a much higher rate using fewer GPUs than prefill, and its optimal parallelism configuration depends on the running batch size. In existing systems, due to coupling, resource allocation and parallelism plans are tailored to satisfy the more demanding of TTFT and TPOT, which may not be ideal for the other. This often leads to resource over-provisioning to meet both SLOs.
@@ 3.1 Analysis for Prefill Instance
Figure 3: Throughput for prefill phase and decoding phase with different batch size and input length when serving an LLM with 13B parameters.
After disaggregation, the prefill phase generates the first token by processing all tokens of the user prompt in parallel. Assuming a given arrival rate, our goal is to fulfill the service’s latency requirement on TTFT using the least resources.
Batching strategy. The prefill step is typically compute-intensive. Figure 3(a) shows how the throughput of the prefill phase changes with the input length and the batch size. For a 13B parameter LLM, processing a single sequence of 512 tokens can fully engage an A100 GPU; larger models require shorter sequences to reach GPU saturation. Once the GPU becomes compute-bound, adding more requests to the batch no longer improves GPU efficiency. Instead, it proportionally extends the total processing time for the batch, inadvertently delaying all included requests. Hence, for prefill instances, it is necessary to profile the specific LLM and GPUs in advance to identify a critical input length threshold, denoted as $L_{m}$, beyond which the prefill phase becomes compute-bound. Batching more requests should only be considered when the input length of the scheduled request is below $L_{m}$. In practice, user prompts typically average over hundreds of tokens [7]. Batch sizes for the prefill instance are generally kept small.
(a) Real experiment results     (b) Changing intra-op speedup
@@
$\small Avg\_TTFT=D+\frac{RD^{2}}{2(1-RD)},$ (1)
where the first term represents the execution time and the second corresponds to the queuing delay. Based on Eq. 1, we incorporate parallelism.
@@
The prefill phase’s preference for parallelism is also influenced by TTFT SLO and the speedup coefficient $K$. Seen from Figure 4(a): A more stringent SLO will make intra-op parallelism more advantageous, due to its ability to support higher request rates while adhering to SLOs. The value of K depends on factors such as the input length, model architecture, communication bandwidth, and placement [39, 50]. As shown in Figure 4(b), a decrease in K notably reduces the efficacy of intra-op parallelism. §4 develops algorithms that optimize the resource and parallelism configurations taking into consideration these knobs.
@@ 3.2 Analysis for Decoding Instance
Unlike the prefill instance, a decoding instance follows a distinct computational pattern: it receives the intermediate states (KV caches) and the first token from the prefill instance and generates subsequent tokens one at a time. For decoding instances, our optimization goal is to satisfy the application’s TPOT requirement using minimal computing resources.
Batching strategy. Since a single decoding job is heavily bandwidth-bound, batching is key to avoiding low GPU utilization (hence high per-gpu goodput). In existing systems where the prefill and decoding phases are colocated, increasing the decoding batch size is difficult because it conflicts with meeting latency goals, particularly in scenarios with high request rates. This is because sharing GPUs cause competition between prefill and decoding jobs, leading to a trade-off between TTFT and TPOT. For example, a higher arrival rate generates more prefill jobs, demanding greater GPU time to meet TTFT requirements if prioritizing prefill jobs, which in turn adversely affects TPOT.
@@
Parallelism plan. Post-disaggregation, the batch size for decoding may be constrained by GPU memory capacity, as it is necessary to maintain the KV caches for all active requests. Scaling the decoding instance with model parallelism or leveraging advanced memory management techniques for LLM KV caches, such as Paged-Attention [28] and GQA [9], enable further scaling the decoding batch size to nearly compute-bound. As the decoding batch size continue to increase to approach the compute bound, the decoding computation begins to resemble the prefill phase. With this observation, we investigate how the latency and throughput change under different parallelism degrees under large batch conditions in Figure 5: intra-op parallelism reduces latency with diminishing returns, caused by communication and reduced utilization after partitioning. Inter-op parallelism can almost linearly scale the throughput. Hence, when the TPOT SLO is stringent, intra-op parallelism is essential to reduce TPOT to meet latency goals. Beyond this, inter-op parallelism is preferable to enhance throughput linearly.
@@ 3.3 Practical Problems
Variable prefill length. §3 has assumed uniform prompt length across requests. In real deployments, depending on the LLM application, the lengths of requests are non-uniform. The non-uniformity can cause pipeline bubbles [25, 31] for prefill instances applying inter-op parallelism, because the execution time of pipeline stages across requests of different lengths will vary. This results in slight deviations from the conclusions indicated by using M/D/1 queue model. To address the problem, §4 develops algorithms that search for parallelisms based on workloads, and resort to scheduling to minimize the bubbles (§4.3).
Communication overhead. Transferring KV caches from prefill to decoding instances incurs notable overheads. For example, the KV cache size of a single 512-token request on OPT-66B is approximately 1.13GB. Assuming an average arrival rate of 10 requests per second, we need to transfer $1.13\times 10=11.3$ GB data – or equivalently 90Gbps bandwidth to render the overhead invisible. The size of the KV caches increases with average input length and arrival rate. While many modern GPU clusters for LLMs are equipped with Infiniband (e.g., 800 Gbps), in cases where cross-node bandwidth is limited, disaggregation relies on the commonly available intra-node NVLINK, where the peak bandwidth between A100 GPUs is 600 GB/s, again rendering the transmission overhead negligible (see §6.3). However, this requirement imposes additional constraints on the placement of preill and decoding instances that we take into consideration in the next section.
@@ 4.1 Placement for High Node-Affinity Cluster
[…]
On high node-affinity clusters equipped with Infiniband, KV caches transmission overhead across nodes is negligible, DistServe can efficiently deploy prefill and decoding instances across any two nodes without constraints. We propose a two-level placement algorithm for such scenarios: we first optimize the parallelism configurations for prefill and decoding instances separately to attain phase-level optimal per-gpu goodput; then, we use replication to match the overall traffic rate.
However, finding the optimal parallel configuration for a single instance type, such as for the prefill instance, is still challenging, due to the lack of a simple analytical formula to calculate the SLO attainment (a.k.a., percentage of requests that meet TTFT requirement), given that the workload has diverse input, output lengths, and irregular arrival patterns. Gauging the SLO via real-testbed profiling is time-prohibitive. We thus resort to building a simulator to estimate the SLO attainment, assuming prior knowledge of the workload’s arrival process and input and output length distributions. Although short-term interval is impossible to predict, the workload pattern over longer timescales (e.g., hours or days) is often predictable [29, 46]. DistServe fits a distribution from the history request traces and resamples new traces from the distribution as the input workload to the simulator to compute the SLO attainment. Next, DistServe simply enumerates the placements via binary search and finds the maximum rate that meets the SLO attainment target with simulation trials.
Algorithm 1 outlines the process. We enumerate all feasible parallel configurations, subject to cluster capacity limit, for both prefill and decoding instances. For example, for a specific prefill phase configuration, we use simu_prefill to simulate and find their maximum goodput (similarly for using simu_decode for decoding). After determining the optimal parallel configurations for both prefill and decoding instances, we replicate them to achieve the user-required overall traffic rate according to their goodput.
@@
By far, we have developed Algorithm 1 assuming we can place the prefill and decoding between any two nodes of the cluster, and the KV cache transmission utilizes high bandwidth. In many real clusters, GPUs inside a node access to high-bandwidth NVLINK while GPUs distributed across nodes have limited bandwidth. We next develop an algorithm to address this constraint.
@@
     $\mathcal{P}\leftarrow\text{get\_intra\_node\_configs}(\hat{G},M,C)$
     for $P\in\mathcal{P}$ do

### v3

# DistServe: Disaggregating Prefill and Decoding for Goodput-optimized Large Language Model Serving

@@ Abstract
DistServe assigns prefill and decoding computation to different GPUs, hence eliminating prefill-decoding interferences. Given the application’s TTFT and TPOT requirements, DistServe co-optimizes the resource allocation and parallelism strategy tailored for each phase. DistServe also places the two phases according to the serving cluster’s bandwidth to minimize the communication caused by disaggregation. As a result, DistServe significantly improves LLM serving performance in terms of the maximum rate that can be served within both TTFT and TPOT constraints on each GPU. Our evaluations show that on various popular LLMs, applications, and latency requirements, DistServe can serve 7.4$\times$ more requests or 12.6$\times$ tighter SLO, compared to state-of-the-art systems, while staying within latency constraints for $>90\%$ of requests.
@@ 1 Introduction
Based on the above insights, in this work, we build DistServe  [^2] https://github.com/LLMServe/DistServe, a goodput-optimized LLM serving system by disaggregating the prefill and decoding phases. Given TTFT and TPOT requirements, DistServe first scales each phase independently by co-optimizing the GPU allocation and parallelism strategies of the prefill and decoding phase assuming serving a single model replica. The optimization ensures maximizing the per-GPU goodput and may assign different numbers of GPUs and parallelism strategies to each phase depending on their respective latency requirements. DistServe then scales this allocation to multiple instances via replication until meeting the user-required traffic rate (§4). DistServe also features an algorithm to place the prefill and decoding computation according to their allocation schemes and the cluster’s bandwidth to minimize the overhead of communicating intermediate states between phases.
We implement DistServe as an orchestration layer on top of the LLM inference engine. We evaluate DistServe on various LLMs, varying the workloads based on three important real-world LLM applications: chatbots, programming assistant, and document summary. Compared to state-of-the-art solutions, DistServe can serve up to 7.4×7.4$\times$7.4 × more requests or 12.6×12.6$\times$12.6 × tighter SLO under various latency constraints. Our contributions are:
  • Identify the problems of prefill-decoding interference and resource coupling in existing LLM serving systems and propose to disaggregate the two phases.
  • Design a novel placement algorithm to choose the goodput-optimal schema for prefill and decoding instances automatically.
@@ 2.1 LLM Inference
The prefill step deals with a new sequence, often comprising many tokens, and processes these tokens concurrently. Unlike prefill, each decoding step only processes one new token generated by the previous step. This leads to significant computational differences between the two phases. When dealing with user prompts that are not brief, the prefill step tends to be compute-bound. For instance, for a 13B LLM, computing the prefill of a 512-token sequence makes an A100 near compute-bound (see §3.1). In contrast, despite processing only one new token per step, the decoding phase incurs a similar level of I/O to the prefill phase, making it constrained by the GPU’s memory bandwidth.
@@ 2.2 LLM Serving Optimization
Batching. Current serving systems [54, 32, 9] utilize a batching technique known as continuous batching. This method batches the prefill of new requests with the decoding of ongoing ones. It boosts the GPU utilization and maximizes the overall system throughput – tokens generated per second across all users and requests. However, as mentioned in §1 and elaborated later in §2.3, this approach leads to trade-offs between TTFT and TPOT. An advanced variant of continuous batching [9] attempts to balance TTFT and TPOT by segmenting long prefill into chunks and attaching decoding jobs with a chunked prefill – but essentially, it trades TTFT for TPOT and cannot eliminate the interference (§2.3). In summary, batching prefill and decoding invariably leads to compromises in either TTFT or TPOT.
Model parallelism. In LLM serving, model parallelism is generally divided as intra- and inter-operator parallelisms [33, 59, 46]. Both can be used to support larger models but may impact serving performance differently. Intra-operator parallelism partitions computationally intensive operators, such as matrix multiplications, across multiple GPUs, accelerating computation but causing substantial communication. It reduces the execution time333we emphasize “execution time” instead of latency here because latency comprises both execution time and queuing delay., hence latency, particularly for TTFT of the prefill phase, but requires high bandwidth connectivity between GPUs (e.g., NVLINK). Inter-operator parallelism organizes LLM layers into stages, each running on a GPU to form pipelines. It moderately increases execution time due to inter-stage communication, but linearly scales the system’s rate capacity with each added GPU. In this paper, we reveal an additional benefit of model parallelism: reduced queuing delay of both prefill and decoding phases, steaming from shorter execution time. We delve into this further in §3. Besides model parallelism, replicating a model instance, irrespective of its model parallelism configurations, linearly scales the system’s rate capacity.
@@ 2.3 Problems and Opportunities
(a) Input length = 128            (b) Input length = 1024
Figure 2: Batch execution time when serving a 13B LLM as batch size increases. Compared between a decoding-only batch and the batch adding one more prefill job.
@@
One attempt to mitigate this interference is called chunked-prefill with piggyback [9, 3]. It proposes to split the long prefill into chunks and batch a prefill chunk with a few decoding jobs (a.k.a. piggybacking). This technique alleviates the slowdown of the decoding job caused by the long prefill job, but it does not eliminate it. Additionally, it results in an extra overhead for the prefill job which cannot be easily mitigated by adjusting the chunk size. First, if the chunk size is set much lower than the inflection point that can saturate the GPU, then the prefill job will have a longer execution time since it competes with the decoding job in the same batch and cannot solely utilize the GPU resources. Second, if we increase the chunk size to nearly saturate the GPU, the chance of piggybacking will diminish since the remaining slots for decode tokens are limited. Also, chunked-prefill causes significantly more memory access for the prefill jobs. This is because the KV cache of all previous chunks have to be loaded from HBM to SRAM repeatedly to compute each subsequent chunk. Concretely, if a prefill job is split into $N$ equal chunks, we need to load $N+(N-1)+...+1=O(N^{2})$ chunks of KV Cache in total, compared to O(N) in the non-chunked case. This overhead will increase as the context length becomes longer.
@@
Resource and parallelism coupling. Colocating prefill and decoding phases on the same GPUs unavoidably share their resource and parallelism settings. However, each phase has its unique computational characteristic and latency requirement that calls for more heterogeneous resource allocation. For example, the prefill phase tends to be compute-bound and benefits from more intra-op parallelism to reduce execution time to meet the tight SLO on TTFT. By contrast, the optimal parallelism configuration of the decoding phase depends on the running batch size. In existing systems, due to coupling, resource allocation and parallelism plans are tailored to satisfy the more demanding of TTFT and TPOT, which may not be ideal for the other. This often leads to resource over-provisioning to meet both SLOs.
@@ 3.1 Analysis for Prefill Instance
Figure 3: Throughput for two phases with different batch sizes and input lengths when serving an LLM with 13B parameters.
After disaggregation, the prefill phase generates the first token by processing all tokens of the user prompt in parallel. Assuming a given arrival rate, we aim to fulfill the service’s latency requirement on TTFT using the least resources.
Batching strategy. The prefill step is typically compute-intensive. Figure 3(a) shows how the throughput of the prefill phase changes with the input length and the batch size. For a 13B parameter LLM, processing a single sequence of 512 tokens can fully engage an A100 GPU. Once the GPU becomes compute-bound, adding more requests to the batch no longer improves GPU efficiency. Instead, it proportionally extends the total processing time for the batch, inadvertently delaying all included requests. Hence, for prefill instances, it is necessary to profile the specific LLM and GPUs in advance to identify a critical input length threshold, denoted as $L_{m}$, beyond which the prefill phase becomes compute-bound. Batching more requests should only be considered when the input length of the scheduled request is below $L_{m}$. In practice, user prompts typically average over hundreds of tokens [8]. Batch sizes for the prefill instance are generally kept small.
@@
$\small Avg\_TTFT=D+\frac{RD^{2}}{2(1-RD)}.$ (1)
where the first term represents the execution time and the second corresponds to the queuing delay. Based on Eq. 1, we incorporate parallelism below.
@@
The prefill phase’s preference for parallelism is also influenced by TTFT SLO and the speedup coefficient $K$. Seen from Figure 4(a): A more stringent SLO will make intra-op parallelism more advantageous, due to its ability to reduce execution time. The value of K depends on factors such as the input length, model architecture, communication bandwidth, and placement [46, 59]. As shown in Figure 4(b), a decrease in K notably reduces the efficacy of intra-op parallelism. §4 develops algorithms that optimize the resource and parallelism configurations taking into consideration these knobs.
(a) Real experiment results     (b) Changing intra-op speedup
@@ 3.2 Analysis for Decoding Instance
Unlike the prefill instance, a decoding instance follows a distinct computational pattern: it receives the KV caches and the first output token from the prefill instance and generates subsequent tokens one at a time. For decoding instances, our optimization goal is to satisfy the application’s TPOT requirement using minimal computing resources.
Batching strategy. Since a single decoding job is heavily bandwidth-bound, batching is key to avoiding low GPU utilization (hence high per-gpu goodput), as shown in Figure 3(b). In existing systems where the prefill and decoding phases are colocated, increasing the decoding batch size is difficult because it conflicts with meeting latency goals, particularly in scenarios with high request rates. This is because sharing GPUs cause competition between prefill and decoding jobs, leading to a trade-off between TTFT and TPOT. For example, a higher arrival rate generates more prefill jobs, demanding greater GPU time to meet TTFT requirements if prioritizing prefill jobs, which in turn adversely affects TPOT.
@@
Parallelism plan. Post-disaggregation, the batch size for decoding may be constrained by GPU memory capacity, as it is necessary to maintain the KV caches for all active requests. Scaling the decoding instance with model parallelism or leveraging advanced memory management techniques for LLM KV caches, such as Paged-Attention [32] and GQA [10], enable further scaling of the decoding batch size to nearly compute-bound. As the decoding batch size continue to increase to approach the compute-bound, the decoding computation begins to resemble the prefill phase. With this observation, we investigate how the latency and throughput change under different parallelism degrees under large batch conditions in Figure 5: intra-op parallelism reduces latency with diminishing returns, caused by communication and reduced utilization after partitioning. Inter-op parallelism can almost linearly scale the throughput. Hence, when the TPOT SLO is stringent, intra-op parallelism is essential to reduce TPOT to meet latency goals. Beyond this, inter-op parallelism is preferable to enhance throughput linearly.
@@ 3.3 Practical Problems
Variable prefill length. §3 has assumed uniform prompt length across requests. In real deployments, depending on the LLM application, the lengths of requests are non-uniform. The non-uniformity can cause pipeline bubbles [28, 36] for prefill instances applying inter-op parallelism because the execution time of pipeline stages across requests of different lengths will vary. This results in slight deviations from the conclusions indicated by using the M/D/1 queue model. To address the problem, §4 develops algorithms that search for parallelisms based on workloads, and resort to scheduling to minimize the bubbles (§4.3).
Communication overhead. Transferring KV caches from prefill to decoding instances incurs notable overheads. For example, the KV cache size of a single 512-token request on OPT-66B is approximately 1.13GB. Assuming an average arrival rate of 10 rps, we need to transfer 11.3GB data per second—or equivalently 90Gbps bandwidth to render the overhead invisible. While many modern GPU clusters for LLMs are equipped with InfiniBand (e.g., 800 Gbps), in cases where cross-node bandwidth is limited, DistServe relies on the commonly available intra-node NVLINK, where the peak bandwidth between A100 GPUs is 600 GB/s, again rendering the transmission overhead negligible (see §6.3). However, this requirement imposes additional constraints on the placement of prefill and decoding instances that we take into consideration in the next section.
@@ 4.1 Placement for High Node-Affinity Cluster
On high node-affinity clusters equipped with Infiniband, KV caches transmission overhead across nodes is negligible, DistServe can deploy prefill and decoding instances across any two nodes without constraints. We propose a two-level placement algorithm for such scenarios: we first optimize the parallelism configurations for prefill and decoding instances separately to attain phase-level optimal per-gpu goodput; then, we use replication to match the overall traffic rate.
However, finding the optimal parallel configuration for a single instance type, such as for the prefill instance, is still challenging, due to the lack of a simple analytical formula to calculate the SLO attainment (a.k.a., percentage of requests that meet TTFT requirement), given that the workload has diverse input, output lengths, and irregular arrival patterns. Gauging the SLO via real-testbed profiling is time-prohibitive. We thus resort to building a simulator to estimate the SLO attainment, assuming prior knowledge of the workload’s arrival process and input and output length distributions. Although short-term interval is impossible to predict, the workload pattern over longer timescales (e.g., hours or days) is often predictable [33, 55]. DistServe fits a distribution from the history request traces and resamples new traces from the distribution as the input workload to the simulator to compute the SLO attainment. Next, DistServe simply enumerates the placements and finds the maximum rate that meets the SLO attainment target with binary search and simulation trials.
Algorithm 1 outlines the process. We enumerate all feasible parallel configurations, subject to cluster capacity limit, for both prefill and decoding instances. Then, for a specific prefill phase configuration, we use simu_prefill to simulate and find its maximum goodput via binary search (similarly for using simu_decode for decoding). After determining the optimal parallel configurations for both prefill and decoding instances, we replicate them to achieve the user-required overall traffic rate according to their goodput.
@@
By far, we have developed Algorithm 1 assuming we can place the prefill and decoding instance between any two nodes (or on the same node) of the cluster, and the KV cache transmission utilizes high bandwidth network. In many real clusters, GPUs inside a node access to high-bandwidth NVLINK while GPUs distributed across nodes have limited bandwidth. We next develop an algorithm to address this constraint.
@@
     for $P_{p}\in\mathcal{P}$ do
         for $P_{d}\in\mathcal{P}$ do
[…]